OpenStudy (anonymous):

Show all work: An excursion takes 1.6 hours longer to go 36 miles up a river than to return. If the rate of the current is 4 miles per hour, what is the rate of the boat in still water?

OpenStudy (anonymous):

help. what do i solve for?

OpenStudy (anonymous):

i think i have the answer can someone just post their result? if i get it wrong then i will need help

OpenStudy (radar):

The problem is asking you to solve for the rate (or speed) of the boat in still water. The problems provides you with the distance (36 miles) It also provides you with the river current rate (4 miles per hour) You are also told that going up river (against the current) takes longer (1.6 hours longer)

OpenStudy (anonymous):

correct... did you get an answer?

OpenStudy (radar):

Not yet, but I let x = speed in still water x+4 = speed downstream x-4 = speed up stream. using that for solving for t with t being t+1.6 for upstream time

OpenStudy (radar):

What did you get for an answer?

OpenStudy (radar):

tu for up stream = 36/(x-4) td for down stream = 36/(x+4) tu-td=1.6 That means: 36/(x-4) - 36/x+4) = 1.6 go ahead and solve for x and see if it agrees with your answer